### v1

# Going Forward-Forward in Distributed Deep Learning

Abstract
This paper introduces a new approach in distributed deep learning, utilizing Geoffrey Hinton’s Forward-Forward (FF) algorithm to enhance the training of neural networks in distributed computing environments. Unlike traditional methods that rely on forward and backward passes, the FF algorithm employs a dual forward pass strategy, significantly diverging from the conventional backpropagation process. This novel method aligns more closely with the human brain’s processing mechanisms, potentially offering a more efficient and biologically plausible approach to neural network training. Our research explores the implementation of the FF algorithm in distributed settings, focusing on its capability to facilitate parallel training of neural network layers. This parallelism aims to reduce training times and resource consumption, thereby addressing some of the inherent challenges in current distributed deep learning systems. By analyzing the effectiveness of the FF algorithm in distributed computing, we aim to demonstrate its potential as a transformative tool in distributed deep learning systems, offering improvements in training efficiency. The integration of the FF algorithm into distributed deep learning represents a significant step forward in the field, potentially revolutionizing the way neural networks are trained in distributed environments.
1 Introduction
Training deep neural networks, often consisting of over hundreds of layers, is a very time consuming process that can take several weeks, using the well-known backpropagation [1] as the standard learning algorithm. While this process works well in a single computational unit, parallelization of these models presents significant challenges due to the sequential nature of the backpropagation learning algorithm, as illustrated in Figure 1. In backpropagation, the gradient of each layer depends on the gradient calculated at the subsequent layer. In a distributed setting, this sequential dependency necessitates extensive waiting times between nodes, as each node must wait for gradient information from its successor before it can proceed with its calculations. Moreover, the need for constant communication between nodes to transfer gradient and weight information can lead to significant communication overhead. This is particularly problematic in large-scale neural networks where the volume of data to be transferred can be substantial.
The realm of distributed deep learning has witnessed significant advancements in recent years, driven by the ever-increasing complexity and size of the neural networks. Distributed training frameworks like GPipe [2], PipeDream [3] and Flower [4] have emerged as pivotal solutions, enabling the training of massive models by optimizing for scale, speed, cost, and usability. These systems utilize sophisticated methodologies such as data, pipeline and model parallelism to manage and process large-scale neural network training efficiently across multiple computing nodes.
Figure 1: Challenges in parallelizing the traditional backpropagation model. F denotes the forward pass, first parameter of F is the layer that the pass is made, and the second parameter is the data that is passed through that layer. B denotes the backward pass and takes the prediction and the real label for calculating the gradient.
Aside from the above research geared towards distributed implementations of backpropagation, the Forward-Forward (FF) algorithm proposed by Hinton [5] proposes a novel approach for training neural networks. Unlike traditional deep learning algorithms that rely on global forward and backward passes, the Forward-Forward algorithm only uses local (layer-wise) computations. While FF algorithm is designed to address the limitations associated with backpropagation without the focus of distributing, layer-wise training feature of FF results in a less dependent architecture in distributed setting that reduces idle time, communication and synchronization as demonstrated in Figure 2.
Figure 2: Parallelization using the FF algorithm. G is the Goodness (to be defined) of the Forward pass.
This paper explores the integration of the FF algorithm into distributed deep learning systems, particularly focusing on its potential to enhance the efficiency and effectiveness of training large-scale neural networks. We demonstrate the effectiveness of the Pipeline Forward-Forward (PFF) algorithm, which reduces the training duration by almost 5-fold for a 5-layer network while preserving the accuracy of the original algorithm.
2 Literature Review
Distributed and parallel neural networks have attracted widespread interest and investigation from researchers, driven by the promise of enhanced computational efficiency and scalability in handling large-scale data and complex models. We used several comprehensive surveys from 2019 [6] and 2022 [7]. These surveys explain the parallelism methods including data, model and pipeline that we employ in our methods.
PipeDream [3] is a deep neural network training system using backpropagation that efficiently combines data, pipeline, and model parallelism by dividing the layers of the model into multiple stages that contains multiple consecutive layers. Unlike our method, PipeDream inherits the limitations of the classic back propagation algorithm.
Similar to PipeDream, GPipe [2] is a distributed deep learning library that updates the weights synchronously during backward pass, which differentiates it from PipeDream. However, it also suffers from additional dependencies introduced by the classic back propagation algorithm.
Forward-Forward algorithm [5] is a learning method for multi-layer networks, proposed by Hinton. This algorithm introduces a novel way of training, which eliminates the back-propagation and backward pass. Different from our method, Forward-Forward Algorithm is sequential, and is designed to work on a single processing unit. This algorithm forms the backbone of our method along with the split method, which aims to reduce communication between workers and increase parallelism. The Forward-Forward algorithm has impacted the field and led to several subsequent works, including neural systems combining Forward-Forward and Predictive Coding [8] and training CNNs with closer accuracy (99.20 accuracy better then the Original FF [5]) to backpropagation using FF. Furthermore, CaFo [9] which is an algorithm that uses FF Algorithm to create multiple cascaded neural blocks which consists of several different layers. In their paper, they also state that the training of different blocks is independent from each other and suitable for parallelization.
Distributed Forward-Forward (DFF) [10] is an example of the use of FF algorithm in distributed training system. DFF aims to create a decentralized and distributed training system to allow multiple low-performance devices to train a large model. DFF creates a cluster with one master node and multiple server nodes. Each server is assigned layer(s) by the master node. These server nodes train their assigned layers using FF Algorithm and send their output to other layers.
While PipeDream and GPipe aim to parallelize backpropagation, DFF [10] is the closest to our work as it also tries to integrate the FF algorithm into a distributed setting. However, in terms of accuracy, our implementation is far superior than DFF because, our implementation updates the weights more frequently, uses minibatches, generates the negative samples adaptively, and uses an additional classifier which all contribute to the performance of the network.
3 Forward-Forward Algorithm
The FF algorithm, proposed by Geoffrey Hinton [5], presents a novel approach to training neural networks. Inspired by the biological processes of the brain, the FF algorithm eliminates backpropagation and utilizes two forward passes to update network weights. Those passes are called respectively, the positive pass and the negative pass.
The positive pass strengthens the recognition of actual data patterns, while the negative pass weakens the recognition for patterns do not match the target data. While other approaches are possible, we will follow Hinton‘s approach where a positive sample is generated by adding the correct label on the top-left corner of a sample image from the training dataset, whereas a negative sample is one where a wrong label is selected (one that does not match the image class).
The most important feature of the FF algorithm for this paper is that its training process doesn’t involve a global backward pass, eliminating the need for backpropagation, which is not deemed as biologically plausible and carries challenges for training deep networks.
Training.
The aim of the learning is to make the ”goodness value” of real (positive) data higher than some threshold and lower it for negative data. The goodness of a layer is simply the sum of the squares of the activities of the rectified linear neurons in that layer. This is then related to the probability that an input vector is positive (i.e., real):
$p(real)=\sigma(\Sigma_{j}y_{j}^{2}-\theta)$
Prediction.
During inference, ’goodness’ score is aggregated across all but the first hidden layer of the network for each label. Then, the label resulting in the maximum goodness is selected as the predicted class.
An alternative inference method was suggested where the input image is labelled with a neutral label (since we do not know the correct class) and a Softmax layer is trained using the activity vectors collected from all layers except the first hidden layer [5].
Figure 3: Traditional Forward-Forward algorithm with three layers and three splits (S = 3).
4 PipelineFF: A Pipeline Forward-Forward Algorithm
In this section we propose three Pipeline Forward Forward (PFF) algorithms: Single Layer PFF (4.1), All Layers PFF (4.2), and Federated PFF (4.3).
The Single Layer PFF employs an architecture where each node is dedicated to training only one layer, while All Layers and Federated PFF algorithms distribute the training load more evenly, by having every node train all network layers in turn. All models enjoy model and pipeline parallelism, while the federated PPF also takes advantage of data parallelism.
In all of the proposed distributed FF models, we split the total number of epochs $E$ into the total number of the splits or chapters $S$ to facilitate distributed training. Each chapter thus takes $C=E/S$ training epochs. Consequently, each client will process $S$ chapters to complete the training for $E$ epochs. The variable $N$ denotes the number of nodes in the distributed system. The input dataset is abstracted as $x$, which is a concatenated form of the positive and negative datasets, simplifying the explanation. $L_{i}(x)$ is used to denote the output of the $i$-th layer of the neural network after a forward pass with the input data $x$. Finally, $Train(L_{i},x,n)$ is an abstracted function that trains layer $L_{i}$ using the data $x$ for $n$ epochs.
4.1 Single Layer PFF Algorithm
Figure 4: Single Layer Pipeline Forward-Forward Algorithm Example with three layers and three splits.
Single Layer PFF employs an architecture where each node is dedicated to training only one layer, as shown in Fig. 4. It reduces data transfer between nodes.
The distributed training example provided in Figure 4 illustrates a scenario with three compute nodes and a three-layer neural network. The number of nodes, $N$, is selected to be equal to the number of layers $L$ and each node trains the corresponding layer of the network. Already with this simple architecture, we achieve parallelism among the nodes, unlike the purely sequential FF algorithm. This procedure ensures that while one node trains a layer, the previous node can simultaneously train the preceding layer. Below, we detail the algorithm with a pseudo code and the training process step-by-step.
Node Initialization. Each node in the distributed computing system, designated as Node $i$ where $i$ ranges from 1 to $N$, is initialized with the necessary data (training data and hyperparameters) to begin training its corresponding layer $L_{i}$.
Initial Training: $Node_{1}$ begins by training the randomly initialized first layer $L_{1}$ using $x$ for $C$ epochs.
Rest of the Layers: Concurrently, as $Node_{i-1}$ completes its $C$ epochs $L_{i-1}$, $Node_{i}$ commences its training of the $L_{i}$, using the output of $L_{i-1}(x)$.
Iteration and Convergence: The process iterates for $S$ chapters, with each node refining its layer based on the progressively improved outputs from the preceding layer. Convergence is achieved when all layers have been trained for $E$ epochs across all splits.
Algorithm 1 Single Layer Training Procedure
1:for each $chapter$ in $range(split)$ do
2:     $x_{pos},x_{neg}\leftarrow X_{POS},X_{NEG}$
3:     for each $layerIndex$ in $range(myIndex)$ do
4:         $layer\leftarrow getLayer(layerIndex,chapter)$
5:         $x_{pos},x_{neg}\leftarrow layer(x_{pos},x_{neg})$
6:     for each $miniEpoch$ in $range(epochs/split)$ do
7:         $learningRateCooldown(chapter,miniEpoch)$
8:         for each $batch$ in $x_{pos},x_{neg}$ do
9:              $trainLayer(myIndex,batch)$
10:     $PublishLayer(chapter,layerIndex)$
11:     $UpdateXNEG(publish=False)$
4.2 All Layers PFF Algorithm
Figure 5: All Layers Pipeline Forward-Forward Algorithm Example with three layers and six splits.
The All Layers PFF model builds upon the Single Layer approach by allowing each compute node to train the entire network, affording a more balanced load among the nodes.
Figure 5 illustrates the approach in the context of a system with three compute nodes and a three-layer neural network, but with six splits for clarity.
Node Initialization. Each node is initialized with the dataset and the required hyper-parameters.
Initial Training: $Node_{1}$ trains the entire network with randomly initialized layers for $C$ epochs.
Rest of the training: Concurrently, each $Node_{i}$ trains layer $L_{k-1}$ immediately after $Node_{i-1}$ has completed its $C$ epochs of training for layer $L_{k-2}$. Each $Node_{i}$ sends its last trained layer to $Node_{i+1(mod(N))}$ and gets the layer to be trained from $Node_{i-1(mod(N))}$.
Iteration and Convergence: The process repeats for $S/N$ but not $S$ since in this approach every node trains each layer. The algorithm converges when the entire network has been trained on total of $E$ epochs.
Algorithm 2 All Layers Training Procedure
1:for each $chapter$ in $range(split)$ do
2:     $x_{pos},x_{neg}\leftarrow X_{POS},X_{NEG}$
3:     for each $layerIndex$ in $range(layerAmount)$ do
4:         if not ($firstClient$ and $chapter=0$then
5:              $getLayer(layerIndex,chapter)$
6:         for each $miniEpoch$ in $range(epochs/split)$ do
7:              $learningRateCooldown(chapter,miniEpoch)$
8:              for each $batch$ in $x_{pos},x_{neg}$ do
9:                  $trainLayer(layerIndex,batch)$
10:         $PublishLayer(chapter,layerIndex)$
11:         $x_{pos},x_{neg}\leftarrow layer(x_{pos},x_{neg})$
12:     $UpdateXNEG(publish=False)$
4.3 Federated PFF Algorithm
In contrast to the All Layers PFF Algorithm, the Federated PFF Algorithm employs the same training procedure but differentiates in the data it utilizes for training. Each node trains on its own local dataset rather than a shared one. This preserves data privacy and autonomy as there is no need to share raw data with other nodes or a centralized server. While the data remains localized, nodes exchange model updates.
This allows nodes to benefit from learning carried out by their counterparts, which iteratively improves the global model that benefits from diverse data without central aggregation.
This federated structure is not only crucial for scenarios requiring data privacy but also enables distributed training over networks where central data aggregation is impractical or undesired.
5 Experimental Evaluation
We obtain and compare several models by modifying the distributed model (Single Layer or All Layers), negative data selection process (Adaptive, Fixed or Random) and classification strategies (Goodness and Softmax).
For selecting negative samples, we implemented Adaptive-NEG, the method proposed in [5], which selects the most predicted incorrect label as the negative label. The second approach is Fixed-NEG, which involves choosing random incorrect labels for each instance in the beginning of the training. The third approach, called Random-NEG, selects the random incorrect labels at the end of each chapter. The last two are evaluated to speed up the training.
As for classifying approach, we adopt the two approaches proposed in [5]. In the Goodness based approach, a sample is classified into the class that results in the highest goodness score (considering all but the first layer). In the Softmax approach, a single layer classification head is added to the network with the softmax activation.
Each of the models are implemented with two PFF variants: Single Layer model (Section  4.1) and All Layers model (Section  4.2). The Sequential PFF, on the other hand, is equivalent to the original FF algorithm, where the node amount $N=1$ for comparing the models using the same code base.
Our chosen architecture for the experiment consists of a neural network with a configuration of [784, 2000, 2000, 2000, 2000] where the numbers indicate the number of nodes in each layer. Thus the input, which is 734-dimensional, is followed by four hidden layers with 2000 nodes each. Each of these layers employs the ReLU activation function and the threshold coefficient is set to 0.01, as in [5]. The network is trained and tested using the MNIST dataset, using 60,000 training instances with mini-batches of size 64 and testing on a separate test set comprising 10,000 instances for 100 epochs and 100 splits. The Adam optimizer is used for both the FF Layers and the Softmax layer (trained using Backpropagation) is set initially at 0.01 and 0.0001 respectively and cooldowns after the 50th epoch.
5.1 Comparison of FF, DFF and PFF Models
We first compare the accuracy and speed of the PFF variants using the Goodness approach, with the results of the original Forward-Forward paper’s Matlab implementation [5] and another distributed implementation [10] of the Forward-Forward algorithm. The results are given in Table 1.
While the results of the compared models are obtained from their respective publications, all our models are trained for the same total number of epochs. The Adaptive-NEG Goodness model with Sequential implementation is basically the same as [5] and it indeed achieves almost the same accuracy as Hinton‘s Matlab implementation [11]. It enables us to compare the training time of the model with our Python implementation.
Among negative data selection methods, the Adaptive-NEG approach performs best, as expected, since it selects the negative data at each chapter. Among the distributive alternatives, the All Layers implementation achieves near top results with over three-fold improvement in training time (2,980 versus 11,190 seconds). Moreover, in the Single Layer method, negative samples in Adaptive-NEG are only generated and published to other nodes from the last node. In contrast, in the All Layers method, each node independently computes the negative samples at the end of each chapter. This way, the tasks are distributed better, concluding to a faster training. Furthermore, to increase time efficiency, in the Single Layer implementation other nodes do not wait for the generated negative samples from the last node at the end of each chapter. Indeed, other nodes begin using the updated negative data immediately upon receiving it. This characteristic slightly reduces the accuracy of the Single Layer model.
Table 1: Original, DFF and PFF Comparison. Baseline is highlighted as bold.
Model Implementation Time (s) Accuracy (%)
Hinton‘s Matlab Code [11] 98.53
DFF (1000 epochs)[10] 93.15
Sequential 11,190.72 98.52
Adaptive-NEG Goodness Single Layer 5,254.87 98.43
All Layers 2,980.76 98.51
Sequential 7,178.71 98.33
Random-NEG Goodness Single Layer 1,974.10 98.26
All Layers 2,008.25 98.17
Sequential 7,143.28 97.95
Fixed-NEG Goodness Single Layer 1,920.80 97.94
All Layers 1,978.21 97.89
Notably, the Random-NEG model achieves a level of accuracy comparable to the Adaptive-NEG model (e.g. 98.17 vs 98.51 accuracy for the All Layers model), while having shorter training times (2008 vs 2980 seconds). Finally, Adaptive-NEG and Random-Neg models performs better than Fixed-NEG model in every implementation, as the variability in the selection of the negative data enhances the robustness of the system.
5.2 Classifier Model Comparison
In the first evaluation, we compared PFF variants that used the Goodness approach for classification and established the Adaptive-NEG Goodness model as the proposed baseline. The second experiment compares the performance of this model with the one using the Softmax approach. As mentioned before, the goodness method is the first method provided in [5] and uses only the FF layers. Softmax method is presented as an alternative, predicting the output using a Softmax layer that gets the activations of the hidden FF layers except the first.
The corresponding results can be found in Table 2. The results show that the Adaptive-NEG Softmax model (results from Table 1) trains faster than the Adaptive-NEG Goodness model, for all different implementations. This accelerated training is due to the single-step prediction in the softmax approach, rather than predicting across 10 different classes. However, since the predictions made by the Softmax layer do not directly indicate the FF Network’s overall performance, we observe a slight drop in accuracy.
Table 2: Softmax Analysis of Adaptive-NEG. The first three row results are from Table 1.
Model Implementation Time (s) Accuracy (%)
Sequential 11,190.72 98.52
Adaptive-NEG Goodness Single Layer 5,254.87 98.43
All Layers 2,980.76 98.51
Sequential 8,365.96 98.38
Adaptive-NEG Softmax Single Layer 2,471.27 98.31
All Layers 1,886.42 98.30
5.3 Softmax Comparison of Random-NEG and Fixed-NEG
In this section, we compare the Softmax method with Fixed-NEG and Random-NEG, with results given in the Table 3.
We first observe that the accuracy of the Softmax approach is slightly higher than the goodness prediction.
As expected, training a model with additional Softmax layer takes longer in the Sequential model. However, Single Layer implementation is not effected from this additional complexity. This is due to the fact that Softmax Layer is easier/faster to train than the FF Layers and thanks to the pipeline architecture this only adds a small delay.
Surprisingly, Softmax models give faster results compared to same Goodness models for All Layers implementation. This again stems from the fact that training Softmax is faster than training FF Layers and distributing this job between nodes speeds ups the overall training.
Table 3: Softmax Comparison of Random-NEG and Fixed-NEG
Model Implementation Time (s) Goodness Softmax
Accuracy (%) Accuracy (%)
Sequential 8,104.96 98.33 98.48
Random-NEG Softmax Single Layer 1,891.86 98.20 98.31
All Layers 1,786.30 98.16 98.33
Sequential 7,178.71 98.33 -
Random-NEG Goodness Single Layer 1,974.15 98.26 -
All Layers 2,008.25 98.17 -
Sequential 7,983.21 97.95 98.03
Fixed-NEG Softmax Single Layer 1,939.62 98.01 97.92
All Layers 1,794.52 97.86 97.85
Sequential 7,143.28 97.95 -
Fixed-NEG Goodness Single Layer 1,920.81 97.94 -
All Layers 1,978.21 97.89 -
6 Conclusion and Future Work
Our work in this paper presents Pipeline Forward-Forward Algorithm (PFF), a novel way of training distributed neural networks using Forward-Forward Algorithm. Compared to the classic implementations with backpropagation and pipeline parallelism [2] [3], PFF is inherently different as it does not enforce the dependencies of backpropagation to the system, thus achieving higher utilization of computational units with less bubbles and idle time. Experiments that are done with PFF shows that PFF Algorithm achieves the same accuracy as standard FF implementation [5] with a 4x speed up. Comparison of PFF with an existing distributed implementation of Forward-Forward (DFF [10]) shows even greater improvements as PFF achieves %5 more accuracy in 10 times less epochs. In addition, the data that is exchanged between layers in PFF is a lot less than DFF since PFF sends the layer information (weights and biases) whereas DFF sends the whole output of the data. This results in less communication overhead compared to DFF.
Besides the exciting results that PFF produced, we believe that our work paves the way for a brand new path in the area of Distributed training of Neural Networks. Thus, there are plenty of different ways that can improve PFF, with some of these approaches outlined below.
  • Exchanging parameters after each batch: In the current implementation of PFF, exchange of parameters between different layers is done after each chapter. Making this exchange after each batch is worth trying since it could tune the weights better and produce higher accuracy. However, it could potentially increase the communication overhead.
  • Usage of PFF in Federated Learning: As PFF trains the model by only exchanging layer information between different nodes, it can be implemented to create a Federated Learning system where each node participates with their own data without sharing it with the other nodes.
  • Different Setups: In the experiments of this work, we used sockets to establish communication between different nodes. This brings additional communication overhead since the data is sent through network. In a setup where computational units of the PFF are closer and can access to a shared resource (Multi GPU architectures) the time of training networks can decrease drastically.
  • Generating Negative Samples Differently: The way that negative samples are generated is a very important aspect of Forward-Forward Algorithm as it directly affects the learning of the network. Thus, discovering new and better ways of generating negative samples would definitely result in a better performing system.
  • Forming an Innovative Framework: A general and efficient framework for training large neural networks can be realized following and improving the novel ideas that are presented in this paper.